OpenStudy (anonymous):

Solve for x: ax+b ----- = 2 cx+d

OpenStudy (zzr0ck3r):

ax+b = 2?

OpenStudy (anonymous):

no this is the equation: \[\frac{ ax+b }{ cx+d } = 2\]

OpenStudy (zzr0ck3r):

multiply both sides by cx+d ax+b = 2(cx+d) distribute the 2 ax+b = 2cx+2d gather the x terms on the left and the rest on the right ax-2cx = 2d-b factor out the x x(a-2c) = 2d-b solve for x x = (2d-b)/(a-2c) note: cx+d cant equal 0 because of our first step.